OpenStudy (anonymous):

What are the x-intercepts of the quadratic function? y = x2 + x – 2 A. 2 and 1 B. 1 C. –2 and 1 D. 2 and –1

OpenStudy (anonymous):

These questions are hard. @Hero @s3a

OpenStudy (s3a):

Factor y = x^2 + x – 2 by asking yourself what equals to the coefficient of x^2 (which is 1) times -2 (the constant term) = -2 and adds up to the coefficient of x (which is also, COINCIDENTALLY, 1). The answer to that is the numbers 1 and -1 such that y = (x+1)(x-1) where x equals either -1 or 1 for y to equal 0. THE VALUES OF x FOR WHICH y = 0 ARE CALLED ZEROES.

OpenStudy (s3a):

Wait, I may have made a mistake.

OpenStudy (s3a):

Yes, I did. Wait.

OpenStudy (s3a):

y = (x-1)(x+2) such that x = 1 or -2

OpenStudy (anonymous):

So its C right?